OpenStudy (anonymous):

What is the y-intercept of the line 3x + 4y = 16?

OpenStudy (anonymous):

-4y=-3x+16 then cancelout the 4 -4/-4y= -3/-4x +16/-4 and y=-3/-4x + -4

OpenStudy (eyust707):

notice that seashell has now put it in slope intercept form!

OpenStudy (eyust707):

that great because we know that form and one of the numbers from it is the intercept!

OpenStudy (eyust707):

y=mx + b

OpenStudy (eyust707):

"b" is the intercept

OpenStudy (anonymous):

thank you.

OpenStudy (eyust707):

no problem!

OpenStudy (eyust707):

what did you get?

OpenStudy (anonymous):

16.